OpenStudy (anonymous):

Fill out the table using the function y=2^(-x) x y -2 -1 0 1 2

OpenStudy (anonymous):

Plug in the following x values and see the output, which will be your y values.

OpenStudy (anonymous):

@sammixboo I'm pretty sure that I got it, but I'm not 100%

sammixboo (sammixboo):

So the x values are known, correct?

OpenStudy (anonymous):

So, show us your y values then.

OpenStudy (anonymous):

Oh wait. I see what they were telling me now. I've got it.

sammixboo (sammixboo):

=)

OpenStudy (anonymous):

So far I've got my y values as x y -2 1/4 -1 1/2 0 1 Does that sound right? @sammixboo

sammixboo (sammixboo):

-2 (wrong) -1 (wrong) 0 (right) For -2, just plug in -2 as x, so you would have y = 2^(--2) which is the same as 2^(2)

sammixboo (sammixboo):

Same goes for -1. 2^(--1) = 2^(1)

OpenStudy (anonymous):

Oh alright. Thank you.

sammixboo (sammixboo):

No problem =)

OpenStudy (anonymous):

After doing that I got: x y -2 -4 -1 -2 0 1 1 2 2 4 This is correct as my final answer, right?.-. (I'm sorry if I'm being bothersome.) @sammixboo

sammixboo (sammixboo):

Not quite. Only the (0, 1) is correct. Let's go through each one, one by one starting with -2

sammixboo (sammixboo):

We know that -2 = x, so plug in -2 for x y = 2^(--2) We know that --2 = 2, so now we have y = 2^(2) 2^2 is the same as 2*2, so what is 2*2?

OpenStudy (anonymous):

4

sammixboo (sammixboo):

Right, so if x = -2 then y = 4 So, for your first one it's just 4 Can you try the second one again?

OpenStudy (anonymous):

It would be just 2 I think.

sammixboo (sammixboo):

Right! So now we have x = -2 then y = 4 x = -1 then y = 2 x = 0 then y = 1 Now, can you do the next one again? (it will be a fraction/decimal)

OpenStudy (anonymous):

 see a pattern?